OpenStudy (anonymous):

[3.06] If g(x) = x2 − 3, find g(4). 13 5 16 −3

OpenStudy (anonymous):

g(4)=4^2 -3

OpenStudy (anonymous):

Thanks man!

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

If you have the time would you be willing to help me with a few more?

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

3.06] If f(x) = x + 2, find the value of x if f(x) = 12. 14 12 2 10